OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \frac{d^{99}y}{dx^{99}}(\sin x)\]

OpenStudy (lgbasallote):

\[y' = \cos x\]

OpenStudy (lgbasallote):

\[y'' = -\sin x\]

OpenStudy (lgbasallote):

\[y''' = -\cos x\]

OpenStudy (lgbasallote):

\[y'''' = \sin x\]

OpenStudy (lgbasallote):

so it repeats every 4 times...

OpenStudy (zarkon):

your original notation is off

OpenStudy (lgbasallote):

lol let's not be concerned about that :))

OpenStudy (zarkon):

lol...ok

OpenStudy (lgbasallote):

99 (mod 4) = 3 so the answer is -cosx?

OpenStudy (zarkon):

yes

OpenStudy (lgbasallote):

wow! nice